Diedre:

Your monthly living expenses are $1500 on an income of $1650 per month. Your goal is to have an emergency fund of 4 times your monthly expenses. Your emergency fund savings account has $2400 and you put the remainder of your monthly income into the emergency fund each month. How much more money would you have to save each month to complete your emergency fund in 12 months? A) $300 per month B) $150 per month C) $75 per month D) $60 per month

dude:

Hey So, this requires a lot of understanding of the parts they gave you `Your monthly living expenses are $1500 on an income of $1650 per month.` Starting with the information they gave you first, you have to spend 1500 out of your 1650 income. [This means that you are left with $150 to save, this means you cannot spend A. $300]


`Your goal is to have an emergency fund of 4 times your monthly expenses.` Now, the next part says you want to save 4 times your expenses (which they mentioned was 1500) [1500\(\times 4\) = 6000] We want to save $6000
`Your emergency fund savings account has $2400 and you put the remainder of your monthly income into the emergency fund each month` You already have 2400 on your saving account. This means we can take this out from the total needed [6000-2400=3600]
`How much more money would you have to save each month to complete your emergency fund in 12 months?` You need to split the amount you need into 12 (12 months time to save) 3600/12 = 300 We can only save half of the amount needed because we only have 150 left This means that we need to save $150 more per month to get to 3600, hence you get B

Diedre:

You are pretty good at this! Thank you very much for your help. I have a few more to go over, so if you want to reply please feel free to do so. I am in an adult wish diploma program and I guess you can say that I am not very good at math. I was far better in English than in this subject. Again thank you for your help.

dude:

Of course, you can make a new post to not overflow this one